# 111$\quad$1

### v1

@@
111$\quad$1
Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, 100190. E-mail: Tong Wu, wutong19s@ict.ac.cn; Yu-Jie Yuan, yuanyujie@ict.ac.cn; Ling-Xiao Zhang, zhanglingxiao@ict.ac.cn; Jie Yang, yangjie01@ict.ac.cn; Lin Gao, gaolin@ict.ac.cn.
222$\quad$2
VAST. Yan-Pei Cao, caoyanpei@gmail.com.
333$\quad$3
Department of Computer Science, University of California. Ling-Qi Yan, lingqi@cs.ucsb.edu.
@@ 1 Introduction
Based on the 3D Gaussian Splatting representation, a number of research works have come out and more are on the way. To help readers get familiar with 3D Gaussian Splatting quickly, we present a survey on 3D Gaussian Splatting, which covers both traditional splatting methods and recent neural-based 3DGS methods. As shown in Fig. 1, we divide these works into three parts by functionality. We first introduce how 3D Gaussian Splatting allows realistic scene reconstruction under various sceneries (Sec. 2) and further present scene editing techniques with 3D Gaussian Splatting (Sec. 3) and how 3D Gaussian Splatting makes downstream applications like digital human possible (Sec. 4). Finally, we summarize recent research works on 3D Gaussian Splatting at a higher level and present future works remaining to be done in this field (Sec. 5). A timeline of representative works can be found in Fig. 2.
@@ 2.1 Point-based Rendering
where $\Sigma^{\prime}$ and $\Sigma=RSS^{T}R^{T}$ are the covariance matrice for 3D Gaussian ellipsoids and projected Gaussian ellipsoids on 2D image from a viewpoint with viewing transformation matrix $W$. $J$ is the Jacobian matrix for the projective transformation. 3DGS shares a similar rendering process with NeRF but there are two major differences between them:
@@ 2.2 Quality Enhancement
Though producing high-quality reconstruction results, there are still improvement spaces for 3DGS’s rendering. Mip-Splatting [31] observed that changing the sampling rate, for example, the focal length, can greatly influence the quality of rendered images by introducing high-frequency Gaussian shape-like artifacts or strong dilation effects. To eliminate the high-frequency Gaussian shape-like artifacts, Mip-Splatting [31] constrains the frequency of the 3D representation to be below half the maximum sampling frequency determined by the training images. In addition, to avoid the dilation effects, it introduces another 2D Mip filter to projected Gaussian ellipsoids to approximate the box filter similar to EWA-Splatting [8]. MS3DGS [32] also aims at solving the aliasing problem in the original 3DGS and introduces a multi-scale Gaussian splatting representation and when rendering a scene at a novel resolution level, it selects Gaussians from different scale levels to produce alias-free images.
@@
Apart from the aliasing problem, the capability of rendering view-dependent effects also needs to be improved. To produce more faithful view-dependent effects, VDGS [33] proposes to model the 3DGS to represent 3D shapes and predict attributes like view-dependent color and opacity with NeRF-like neural network instead of Spherical Harmonic (SH) coefficients in the original 3DGS. Scaffold-GS [34] proposes to initialize a voxel grid and attach learnable features onto each voxel point and all attributes of Gaussians are determined by interpolated features and lightweight neural networks. Instead of changing the view-dependent appearance modeling approach, StopThePop [35] points out that 3DGS tends to cheat view-dependent effects by popping 3D Gaussians due to the per-ray depth sorting, which leads to less faithful results when the viewpoint is rotated. To mitigate the potential of popping 3D Gaussians, StopThePop [35] replaces the per-ray depth sorting with tile-based sorting to ensure consistent sorting order at a local region. To better guide the growth of 3D Gaussian Splatting, GaussianPro [30] introduces a progressive propagation strategy to updating Gaussians by considering the normal consistency between neighboring views and adding plane constraints as shown in Fig. 3. To deal with more complex shadings like specular and anisotropic components, Spec-Gaussian [36] proposes to utilize Anisotropic Spherical Gaussian to approximate 3D scenes’ appearance. Quantitative results of different reconstruction methods can be found in Table 1. 3DGS-based methods and NeRF-based methods are comparable but 3DGS-based methods have faster rendering speed.
@@ 2.3 Compression and Regularization
Vector Quantization, a traditional compression method in signal processing, which involves clustering multi-dimensional data into a finite set of representations, is mainly utilized in Gaussians [40, 41, 42, 43, 44]. C3DGS [40] adopts residual vector quantization (R-VQ) [45] to represent geometric attributes, including scaling and rotation. SASCGS [42] utilizes vector clustering to encode color and geometric attributes into two codebooks, with a sensitivity-aware K-Means method. As shown in Fig. 4, EAGLES [43] quantizes all attributes including color, position, opacity, rotation, and scaling, they show that the quantization of opacity leads to fewer floaters or visual artifacts in the novel view synthesis task. Compact3D [41] does not quantize opacity and position, because sharing them results in overlapping Gaussians. LightGaussian [44] adopts octree-based lossless compression in G-PCC [46] for the position attribute due to the sensitivity to the subsequent rasterization accuracy for the position. SOGS [47] adopt a different method from Vector Quantization. They arrange Gaussian attributes into multiple 2D grids. These grids are sorted and a smoothness regularization is applied to penalize all pixels that have very different values compared to their local neighborhood on the 2D grid.
@@ Methods
[…]
3DGS[5] 0.815 27.21 0.214
750
C3DGS[40]
0.798
27.08
0.247
48.8
Compact3D [41]
0.808
27.16 0.228
EAGLES [43] 0.81
27.15
0.24
68
SOGS[47]
0.763
25.83
0.273
18.2
SASCGS[42]
0.801
26.98
0.238
28.80
@@ 2.4 Dynamic 3D Reconstruction
The same as the NeRF representation, 3DGS can also be extended to reconstruct dynamic scenes. The core of dynamic 3DGS lies in how to model the variations of Gaussian attribute values over time. The most straightforward way is to assign different attribute values to 3D Gaussian at different timesteps. Luiten et al. [16] regard the center and rotation (quaternion) of 3D Gaussian as variables that change over time, while other attributes remain constant over all timesteps, thus achieving 6-DOF tracking by reconstructing dynamic scenes. However, the frame-by-frame discrete definition lacks continuity, which can cause poor results in long-term tracking. Therefore, physical-based constraints are introduced, which are three regularization losses, including short-term local-rigidity and local-rotation similarity losses and a long-term local-isometry loss. However, this method still lacks inter-frame correlation and requires high storage overhead for long-term sequences. Therefore, decomposing spatial and temporal information and modeling them with a canonical space and a deformation field, respectively, has become another exploration direction. The canonical space is the static 3DGS, then the problem becomes how to model the deformation field. One way is to use an MLP network to implicitly fit it, similar to the dynamic NeRF [48]. Yang et al. [17] follow this idea and propose to input the positional-encoded Gaussian position and timestep $t$ to the MLP which outputs the offsets of the position, rotation, and scaling of 3D Gaussian. However, inaccurate poses may affect rendering quality. This is not significant in continuous modeling of NeRF, but discrete 3DGS can amplify this problem, especially in the time interpolation task. So, they add a linearly decaying Gaussian noise to the encoded time vector to improve temporal smoothing without additional computational overhead. Some results are shown in Fig. 5. 4D-GS [49] adopts the multi-resolution HexPlane voxels [50] to encode the temporal and spatial information of each 3D Gaussian rather than positional encoding and utilizes different compact MLPs for different attributes. For stable training, it first optimizes static 3DGS and then optimizes the deformation field represented by an MLP. GauFRe [51] applies the exponential and normalization operation to the scaling and rotation respectively after adding the delta values predicted by an MLP, ensuring convenient and reasonable optimization. As dynamic scenes contain large static parts, it randomly initializes the point cloud into dynamic point clouds and static point clouds, optimizes them accordingly, and renders them together to achieve decoupling of the dynamic part and the static part.
@@
These dynamic modeling methods can be further applied in the medical field, such as the markless motion reconstruction for motion analysis of infants and neonates [64] which introduces additional mask and depth supervisions, and monocular endoscopic reconstruction [65, 66]. Quantitative reconstruction results by representative NeRF-based and 3DGS-based methods are reported in Table 3. 3DGS-based methods have clear advantages compared to NeRF-based methods due to their explicit geometry representation that can model the dynamics more easily.
@@
While most methods experiment on regular input data with dense viewpoints in relatively small scenes, there are also works targeting reconstructing 3D scenes with challenging inputs like sparse-view input, data without camera parameters, and larger scenes like urban streets. FSGS [67] is the first to explore reconstructing 3D scenes from sparse view input. It initializes sparse Gaussians from structure-from-motion (SfM) methods and identifies them by unpooling existing Gaussians. To allow faithful geometry reconstruction, an extra pre-trained 2D depth estimation network helps to supervise the rendered depth images. SparseGS [68] also targets 3D reconstruction from sparse-view inputs by introducing depth inputs estimated by pre-trained a 2D network. It further removes Gaussians with incorrect depth values and utilizes the Score Distillation Sampling (SDS) loss [69] to encourage rendered results from novel viewpoints to be more faithful. GaussainObject [70] instead initializes Gaussians with visual hull and fine-tunes a pre-trained ControlNet [71] repair degraded rendered images generated by adding noises to Gaussians’ attributes, which outperforms previous NeRF-based sparse-view reconstruction methods as shown in Fig. 6.
Moving a step forward, pixelSplat [72] reconstructs 3D scenes from single-view input without any data priors. It extracts pixel-aligned image features similar to PixelNeRF [73] and predicts attributes for each Gaussian with neural networks. SplatterImage [74] also works on single-view data but instead utilizes a U-Net [75] network to translate the input image into attributes on Gaussians. It can extend to multi-view inputs by aggregating predicted Gaussians from different viewpoints via warping operation.
@@
For urban scene data, PVG [78] makes Gaussian’s mean and opacity value time-dependent functions centered at corresponding Gaussian’s life peak (maximum prominence over time). DrivingGaussian [79] reconstructs dynamic driving data by first incrementally optimizing static 3D Gaussians and then composing them with dynamic objects’ 3D Gaussians. This process is also assisted by the Segmentation Anything Model [80] and input LiDAR depth data. StreetGaussians [81] models the static background with a static 3DGS and dynamic objects by a dynamic 3DGS where Gaussians are transformed by tracked vehicle poses and their appearance is approximated with time-related Spherical Harmonics (SH) coefficients. To facilitate comparisons on the urban scenes for 3DGS methods, GauU-Scene [82] provides a large-scale dataset covering over $1.5km^{2}$.
Apart from the works mentioned above, other methods focus on special input data including images without camera [83, 84], blurry inputs [85], CT scans [86], panoramic images [87], and satellite images [88].

### v2

@@
111$\quad$1 Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China, 100190. E-mail: Tong Wu, wutong19s@ict.ac.cn; Yu-Jie Yuan, yuanyujie@ict.ac.cn; Ling-Xiao Zhang, zhanglingxiao@ict.ac.cn; Jie Yang, yangjie01@ict.ac.cn; Lin Gao, gaolin@ict.ac.cn.
222$\quad$2 VAST. Yan-Pei Cao, caoyanpei@gmail.com.
333$\quad$3 Department of Computer Science, University of California. Ling-Qi Yan, lingqi@cs.ucsb.edu.
@@ 1 Introduction
Based on the 3D Gaussian Splatting representation, a number of research works have come out and more are on the way. To help readers get familiar with 3D Gaussian Splatting quickly, we present a survey on 3D Gaussian Splatting, which covers both traditional splatting methods and recent neural-based 3DGS methods. There are two literature reviews [6, 7] that summarize recent works on 3DGS, which also serve as a good reference tool for this field. As shown in Fig. 1, we divide these works into three parts by functionality. We first introduce how 3D Gaussian Splatting allows realistic scene reconstruction under various sceneries (Sec. 2) and further present scene editing techniques with 3D Gaussian Splatting (Sec. 3) and how 3D Gaussian Splatting makes downstream applications like digital human possible (Sec. 4). Finally, we summarize recent research works on 3D Gaussian Splatting at a higher level and present future works remaining to be done in this field (Sec. 5). A timeline of representative works can be found in Fig. 2.
@@ 2.1 Point-based Rendering
where $\Sigma^{\prime}$ and $\Sigma=RSS^{T}R^{T}$ are the covariance matrices for 3D Gaussian ellipsoids and projected Gaussian ellipsoids on 2D image from a viewpoint with viewing transformation matrix $W$. $J$ is the Jacobian matrix for the projective transformation. 3DGS shares a similar rendering process with NeRF but there are two major differences between them:
@@ 2.2 Quality Enhancement
Though producing high-quality reconstruction results, there are still improvement spaces for 3DGS’s rendering. Mip-Splatting [33] observed that changing the sampling rate, for example, the focal length, can greatly influence the quality of rendered images by introducing high-frequency Gaussian shape-like artifacts or strong dilation effects. To eliminate the high-frequency Gaussian shape-like artifacts, Mip-Splatting [33] constrains the frequency of the 3D representation to be below half the maximum sampling frequency determined by the training images. In addition, to avoid the dilation effects, it introduces another 2D Mip filter to projected Gaussian ellipsoids to approximate the box filter similar to EWA-Splatting [10]. MS3DGS [34] also aims at solving the aliasing problem in the original 3DGS and introduces a multi-scale Gaussian splatting representation and when rendering a scene at a novel resolution level, it selects Gaussians from different scale levels to produce alias-free images. Analytic-Splatting [35] approximates the cumulative distribution function of Gaussians with a logistic function to better model each pixel’s intensity response for anti-alias. SA-GS [36] utilizes an adaptive 2D low-pass filter at the test time according to the rendering resolution and camera distance.
@@
Apart from the aliasing problem, the capability of rendering view-dependent effects also needs to be improved. To produce more faithful view-dependent effects, VDGS [37] proposes to model the 3DGS to represent 3D shapes and predict attributes like view-dependent color and opacity with NeRF-like neural network instead of Spherical Harmonic (SH) coefficients in the original 3DGS. Scaffold-GS [38] proposes to initialize a voxel grid and attach learnable features onto each voxel point and all attributes of Gaussians are determined by interpolated features and lightweight neural networks. Based on Scaffold-GS, Octree-GS [39] introduces a level-of-detail strategy to better capture details. Instead of changing the view-dependent appearance modeling approach, StopThePop [40] points out that 3DGS tends to cheat view-dependent effects by popping 3D Gaussians due to the per-ray depth sorting, which leads to less faithful results when the viewpoint is rotated. To mitigate the potential of popping 3D Gaussians, StopThePop [40] replaces the per-ray depth sorting with tile-based sorting to ensure consistent sorting order at a local region. To better guide the growth of 3D Gaussian Splatting, GaussianPro [32] introduces a progressive propagation strategy to updating Gaussians by considering the normal consistency between neighboring views and adding plane constraints as shown in Fig. 3. GeoGaussian [41] proposes to densify Gaussians on the tangent plane of Gaussians and encourage the smoothness of geometric properties between neighboring Gaussians. RadSplat [42] initializes 3DGS with point cloud derived from a trained neural radiance field and prunes Gaussians with a multi-view importance score. To deal with more complex shadings like specular and anisotropic components, Spec-Gaussian [43] proposes to utilize Anisotropic Spherical Gaussian to approximate 3D scenes’ appearance. TRIPS [44] attaches a neural feature onto Gaussians and renders pyramid-like image feature planes according to the projected Gaussians size similar to ADOP [45] to resolve the blur issue in the original 3DGS. Handling the same issue, FreGS [46] applies frequency-domain regularization to the rendered 2D image to encourage high-frequency details recovery. GES [47] utilizes the Generalized Normal Distribution(NFD) to sharpen scene edges. To resolve the problem that 3DGS is sensitive to initialization, RAIN-GS [48] proposes to initialize Gaussians sparsely with large variance from the SfM point cloud and progressively applies low-pass filtering to avoid 2D Gaussian projections smaller than a pixel. Pixel-GS [49] takes the number of pixels that a Gaussian covers from all the input viewpoints into account in the splitting process and scales the gradient according to the distance to the camera to suppress floaters. Bulò et al. [50] also leverage pixel-level error as a densification criteria and revise the opacity setting in the cloning process for more stable training process. Quantitative results of different reconstruction methods can be found in Table 1. 3DGS-based methods and NeRF-based methods are comparable but 3DGS-based methods have faster rendering speed.
@@ 2.3 Compression and Regularization
Vector Quantization, a traditional compression method in signal processing, which involves clustering multi-dimensional data into a finite set of representations, is mainly utilized in Gaussians [54, 55, 56, 57, 58]. C3DGS [54] adopts residual vector quantization (R-VQ) [59] to represent geometric attributes, including scaling and rotation. SASCGS [56] utilizes vector clustering to encode color and geometric attributes into two codebooks, with a sensitivity-aware K-Means method. As shown in Fig. 4, EAGLES [57] quantizes all attributes including color, position, opacity, rotation, and scaling, they show that the quantization of opacity leads to fewer floaters or visual artifacts in the novel view synthesis task. Compact3D [55] does not quantize opacity and position, because sharing them results in overlapping Gaussians. LightGaussian [58] prunes Gaussians with a smaller importance score and adopts octree-based lossless compression in G-PCC [60] for the position attribute due to the sensitivity to the subsequent rasterization accuracy for the position. Based on the same importance score calculation, Mini-Splatting [61] samples Gaussians instead of pruning points to avoid artifacts caused by pruning. SOGS [62] adopt a different method from Vector Quantization. They arrange Gaussian attributes into multiple 2D grids. These grids are sorted and a smoothness regularization is applied to penalize all pixels that have very different values compared to their local neighborhood on the 2D grid. HAC [63] adopts the idea of Scaffold-GS [38] to model the scene with a set of anchor points and learnable features on these anchor points. It future introduces an adaptive quantization module to compress the features of anchor points with the multi-resolution hash grid [2]. Jo et al. [64] propose to identify unnecessary Gaussians to compress 3DGS and accelerate the computation. Apart from 3D compression, the 3D Gaussian Splatting has also been applied to 2D image compression [65], where the 3D Gaussians degenerates to 2D Gaussians.
@@
3DGS[5] 0.815 27.21 0.214 750
C3DGS[54] 0.798 27.08 0.247 48.8
Compact3D [55] 0.808 27.16 0.228 /
EAGLES [57] 0.81 27.15 0.24 68
SOGS[62] 0.763 25.83 0.273 18.2
SASCGS[56] 0.801 26.98 0.238 28.80
@@ 2.4 Dynamic 3D Reconstruction
The same as the NeRF representation, 3DGS can also be extended to reconstruct dynamic scenes. The core of dynamic 3DGS lies in how to model the variations of Gaussian attribute values over time. The most straightforward way is to assign different attribute values to 3D Gaussian at different timesteps. Luiten et al. [18] regard the center and rotation (quaternion) of 3D Gaussian as variables that change over time, while other attributes remain constant over all timesteps, thus achieving 6-DOF tracking by reconstructing dynamic scenes. However, the frame-by-frame discrete definition lacks continuity, which can cause poor results in long-term tracking. Therefore, physical-based constraints are introduced, which are three regularization losses, including short-term local-rigidity and local-rotation similarity losses and a long-term local-isometry loss. However, this method still lacks inter-frame correlation and requires high storage overhead for long-term sequences. Therefore, decomposing spatial and temporal information and modeling them with a canonical space and a deformation field, respectively, has become another exploration direction. The canonical space is the static 3DGS, then the problem becomes how to model the deformation field. One way is to use an MLP network to implicitly fit it, similar to the dynamic NeRF [66]. Yang et al. [19] follow this idea and propose to input the positional-encoded Gaussian position and timestep $t$ to the MLP which outputs the offsets of the position, rotation, and scaling of 3D Gaussian. However, inaccurate poses may affect rendering quality. This is not significant in continuous modeling of NeRF, but discrete 3DGS can amplify this problem, especially in the time interpolation task. So, they add a linearly decaying Gaussian noise to the encoded time vector to improve temporal smoothing without additional computational overhead. Some results are shown in Fig. 5. 4D-GS [67] adopts the multi-resolution HexPlane voxels [68] to encode the temporal and spatial information of each 3D Gaussian rather than positional encoding and utilizes different compact MLPs for different attributes. For stable training, it first optimizes static 3DGS and then optimizes the deformation field represented by an MLP. GauFRe [69] applies the exponential and normalization operation to the scaling and rotation respectively after adding the delta values predicted by an MLP, ensuring convenient and reasonable optimization. As dynamic scenes contain large static parts, it randomly initializes the point cloud into dynamic point clouds and static point clouds, optimizes them accordingly, and renders them together to achieve decoupling of the dynamic part and the static part. 3DGStream [70] allows online training of 3DGS into the dynamic scene reconstruction by modeling the transformation between frames as a neural transformation cache and adaptively adding 3D Gaussians to handle emerging objects. 4DGaussianSplatting [71] turns 3D Gaussians into 4D Gaussians and slice the 4D Gaussians into 3D Gaussians for each time step. The sliced 3D Gaussians are projected onto image plane to reconstruct the corresponding frame. Guo et al. [72] and GaussianFlow [73] introduce 2D flow estimation results into the training of dynamic 3D Gaussians, which supports deformation modeling between neighboring frames and enables superior 4D reconstruction and 4D generation results. TOGS [74] constructs an opacity offset table to model the changes in digital subtraction angiography. Zhang et al. [75] leverage the diffusion prior to enhance dynamic scene reconstruction and propose a neural bone transformation module for animatable objects reconstruction from monocular videos.
@@
These dynamic modeling methods can be further applied in the medical field, such as the markless motion reconstruction for motion analysis of infants and neonates [88] which introduces additional mask and depth supervisions, and monocular endoscopic reconstruction [89, 90, 91, 92]. Quantitative reconstruction results by representative NeRF-based and 3DGS-based methods are reported in Table 3. 3DGS-based methods have clear advantages compared to NeRF-based methods due to their explicit geometry representation that can model the dynamics more easily. The efficient rendering of 3DGS also avoids densely sampling and querying the neural fields in NeRF-based methods and makes downstream applications of dynamic reconstruction like free-viewpoint video more feasible.
@@
While most methods experiment on regular input data with dense viewpoints in relatively small scenes, there are also works targeting reconstructing 3D scenes with challenging inputs like sparse-view input, data without camera parameters, and larger scenes like urban streets. FSGS [93] is the first to explore reconstructing 3D scenes from sparse view input. It initializes sparse Gaussians from structure-from-motion (SfM) methods and identifies them by unpooling existing Gaussians. To allow faithful geometry reconstruction, an extra pre-trained 2D depth estimation network helps to supervise the rendered depth images. SparseGS [94], CoherentGS [95], and DNGaussian [96] also target 3D reconstruction from sparse-view inputs by introducing depth inputs estimated by pre-trained a 2D network. It further removes Gaussians with incorrect depth values and utilizes the Score Distillation Sampling (SDS) loss [97] to encourage rendered results from novel viewpoints to be more faithful. GaussainObject [98] instead initializes Gaussians with visual hull and fine-tunes a pre-trained ControlNet [99] repair degraded rendered images generated by adding noises to Gaussians’ attributes, which outperforms previous NeRF-based sparse-view reconstruction methods as shown in Fig. 6. Moving a step forward, pixelSplat [100] reconstructs 3D scenes from single-view input without any data priors. It extracts pixel-aligned image features similar to PixelNeRF [101] and predicts attributes for each Gaussian with neural networks. MVSplat [102] brings the cost volume representation into sparse view reconstruction, which is taken as input for the attributes prediction network for Gaussians. SplatterImage [103] also works on single-view data but instead utilizes a U-Net [104] network to translate the input image into attributes on Gaussians. It can extend to multi-view inputs by aggregating predicted Gaussians from different viewpoints via warping operation.
@@
For urban scene data, PVG [107] makes Gaussian’s mean and opacity value time-dependent functions centered at corresponding Gaussian’s life peak (maximum prominence over time). DrivingGaussian [108] and HUGS [109] reconstruct dynamic driving data by first incrementally optimizing static 3D Gaussians and then composing them with dynamic objects’ 3D Gaussians. This process is also assisted by the Segmentation Anything Model [110] and input LiDAR depth data. StreetGaussians [111] models the static background with a static 3DGS and dynamic objects by a dynamic 3DGS where Gaussians are transformed by tracked vehicle poses and their appearance is approximated with time-related Spherical Harmonics (SH) coefficients. SGD [112] incorporate diffusion priors with street scene reconstruction to improve the novel view synthesis results similar to ReconFusion [113]. HGS-Mapping [114] separately models textureless sky, ground plane, and other objects for more faithful reconstruction. VastGaussian [115] divides a large scene into multiple regions based on the camera distribution projected on the ground and learns to reconstruct a scene by iteratively adding more viewpoints into training based on visibility criteria. In addition, it models the appearance changes with an optimizable appearance embedding for each view. CityGaussian [116] also models the large-scale scene with a divide-and-conquer strategy and further introduces level-of-detail rendering based on the distance between the camera to a Gaussian. To facilitate comparisons on the urban scenes for 3DGS methods, GauU-Scene [117] provides a large-scale dataset covering over $1.5km^{2}$.
Apart from the works mentioned above, other methods focus on special input data including images without camera [118, 119, 120, 121], blurry inputs [122, 123, 124, 125], unconstrained images [126, 127], mirror-like inputs [128, 129], CT scans [130, 131], panoramic images [132], and satellite images [133].